Group Selection Page HELP Controls Group Data Analysis Group ITDM WP76 DET Vacuum EEE WP78 WP74 FXE SCS SPB MID HED SQS SXP Sample Environment Operation PSPO XO TS Migrated2Zulip
OPERATION-2025 OPERATION-2024 OPERATION-2023 OPERATION-2022 OPERATION-2021 OPERATION-2020 OPERATION-2019 SASE1-2018 SASE1-2017 Experiment data
  OPERATION-2023 logbook, Contact: F.Wolff-Fabris, A.Galler, L.Samoylova, H.Sinn  Not logged in ELOG logo
Message ID: 68     Entry time: 03 Mar 2023, 23:25
Author: Giuseppe, Sergii 
Group: SCS 
Subject: Late shift summary 

achievements:

  • change the optical laser to 2 um
  • found focus of 2 um beam
  • found saptial overlap
  • course timing was checked
  • looking for fine timing using SiN

problems:

  • fine timing signal not seen yet
ELOG V3.1.4-7c3fd00